OpenStudy (anonymous):

how do i solve this bernoulli differential equation?? y' + xy = 6xy^(1/2) i got to the part where you find the linear equation and i got.. z' = 3x + x/2y but I dont know who to solve that equation.

OpenStudy (anonymous):

Google :)

OpenStudy (amistre64):

tell bernoulli to do his own work or else hell never learn :)

OpenStudy (jamesj):

Divide your original equation by y^(1/2) and you have y'/y^(1/2) + x y^(1/2) = .... Now set u = y^(1/2) then u' = (1/2) y'/y^(1/2) Substitute for y and y' in the equation above and now you should be able to solve the equation for u.

OpenStudy (jamesj):

Tracking?